OpenStudy (anonymous):

Solve x2 + 10x + 7 = 0 by completing the square. Which equation is used in the process?

OpenStudy (anonymous):

x^2 + 10x + 7 = 0 x^2 + 10x = -7 10 / 2 = 5 5^2 = 25 x^2 + 10x + 25 = -7 + 25 (x + 5)^2 = 18 x = -5 +/- 3(sqrt)2

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You're welcome! And thanks for the medal. :)